OpenStudy (anonymous):

The function C (n) = 1.25n represents the cost, C(n), when you purchase n sodas. What is the cost of purchasing 12 sodas?

OpenStudy (anonymous):

plug in n = 12

OpenStudy (the_fizicx99):

C(12) = 1.25*12 = ?

OpenStudy (the_fizicx99):

Do you understand? 1.25 * 12 = 15

OpenStudy (anonymous):

Oh yes I understand now.

OpenStudy (the_fizicx99):

Yw :)

OpenStudy (anonymous):

Thank youuuuu :D

OpenStudy (the_fizicx99):

noo medal :(